OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (anonymous):

Hello there! I can help you out. Do you have any idea where to start with this, or are you completely lost?

OpenStudy (anonymous):

im completely lost!

OpenStudy (anonymous):

That's perfectly fine :) The thing about math is you really have to get those formulas down, know what I mean? :) That being said, you should really get the formula for this problem memorized because teachers are going to be drilling you with it for a while! Here's the equation you need for slope-intercept form: y = mx + b where "m" is the slope of the line and "b" is the y-intercept. Does this make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, good. So now you can start plugging in the variables for your problem. y = mx + b Well, they gave you the slope, right? So which variable (letter) would you take out and replace with your slope (3)?

OpenStudy (anonymous):

the m would be 3. so y = 3x + b

OpenStudy (anonymous):

Perfect! Okay, so do you know what your y intercept would be?

OpenStudy (anonymous):

its -2 right?

OpenStudy (anonymous):

Yes! Perfect :)

OpenStudy (anonymous):

-2= 3x + b

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I have one more question. That i need some help on. Could you help me real fast please?

OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

WAIT! I overlooked something very important. I was thinking of another slope formula (see why I said to make sure to keep your formulas in order? *sheepish grin*). Take it back to when you wrote "-2= 3x + b" You were correct in writing it like that. There are a few more steps to take once you reached that point.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2= 3x + b is still following the y = mx + b thing. You've plugged in m(slope), you've plugged in your y-int (-2) and now you have to plug in your x-intercept. What should your equation look like now?

OpenStudy (anonymous):

-2 = 3x + 1?

OpenStudy (anonymous):

No. We're solving for b. (But that was a good guess.) This is what we're looking for: -2x = 3(1) +b Do you understand why the equation looks like this?

OpenStudy (anonymous):

yeah but why did we add an x to the -2?

OpenStudy (anonymous):

Oops, my mistake. Ignore that. :>

OpenStudy (anonymous):

-2 = 3(1) + b

OpenStudy (anonymous):

Now... Could you solve for b?

OpenStudy (anonymous):

would it be 3?

OpenStudy (anonymous):

Hmm, not sure how you got that.... -2 = 3(1) + b [so, multiply 3 * 1] -2 = 3 + b [simplified] Now you have to solve for b, which means you want b to be by itself. So you would have to add -3 to both sides (so you cancel 3 out, and get b alone on one side of the equation). Make sense?

OpenStudy (anonymous):

umm.. not really?

OpenStudy (anonymous):

That's okay! :) I'll show you what I mean. First, we have to cancel out this 3 over here by the b.... -2 = 3 + b <------- we don't want 3 over here. We're solving for b, so this 3 is crowding b's space, popping its bubble. :) We have to move it to the other side of the equation, next to -2, which doesn't mind having a buddy right now. :) (-3 - 2) = (3 - 3) + b <--- We get rid of that 3 by adding it to -3, on BOTH SIDES. On the 'b' side: When you add a negative 3 and a positive 3 together, it makes zero (it cancels out). And you also add -3 to -2, because whatever you do to one side, you do to the other. So your equation will look like this: -5 = 0 + b which is another way of saying: -5 = b Yes?

OpenStudy (anonymous):

Yes! :)

OpenStudy (anonymous):

Great! Now we can finally plug 'b' in our equation from earlier. y = mx + b We got the slope, and now we FINALLY know the y-intercept (b). So our equation should look like?

OpenStudy (anonymous):

y = 3x + 2

OpenStudy (anonymous):

Well, we just found out that "b" equals -5. y = mx + b..... y = 3x - 5 Do you get this?

OpenStudy (anonymous):

ohhhhhhh...... lol i see know!

OpenStudy (anonymous):

Excellent! Here's a list of the steps we went through, so you can see it easier. This might help later when you have other problems like this. y = mx + b, given slope = 3, point (1, -2) 1.) Plug in the slope: y = 3x + b 2.) Plug in the x- and y-intercepts (of the point given): -2 = 3(1) + b 3.) Simplify: -2 = 3 + b 4a.) Get 'b' alone on one side: (-3 - 2) = (3 - 3) + b 4b.) Simplify: -5 = b 5.) Plug 'b' into your equation (this is your final answer): y = 3x - 5 Does this help?

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

You're welcome! So did you need help on that other equation now? If so, could you please make another post? This one is getting rather long.

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

Okay :)